If ​[H3O+] = 10-6M, provide the hydroxide ion concentration ​[HO-] by performing a calculation using equation Kw = [H3O+][HO-] = 1x10-14. Please input your answer in the format 1E-7 (with no units).
Answer #1

Given:
[H3O+] = 1*10^-6 M

use:
[OH-] = Kw/[H3O+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H3O+]
[OH-] = (1.0*10^-14)/(1*10^-6)
[OH-] = 1*10^-8 M
Answer: 1E-8